OpenStudy (anonymous):

If the range of the function f(x) = 4 − 4x is {-20, -16, -8, 0, 4}, what is its domain?

OpenStudy (anonymous):

{0, 1, 2, 4, 6} {2, 3, 5, 6, 7} {1, 3, 5, 7, 9} {0, 1, 3, 5, 6}

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

ok, you HAVE TO work on your problem, not me!! I guide you solve for x: 4-4x =-20 , then x =?? 4-4x= 4 , then x =??

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

Do one by one, ok? 4 -4x =-20 -4 both sides, what do you get?

OpenStudy (anonymous):

-4x = -24

OpenStudy (loser66):

divided both sides by -4, what do you get?

OpenStudy (anonymous):

x=6

OpenStudy (loser66):

yup, let it there. x =6 now, next, do the same with 4-4x =-16, what do you get for x?

OpenStudy (anonymous):

-4x = -20

OpenStudy (loser66):

x =?

OpenStudy (anonymous):

x= 5

OpenStudy (loser66):

so far, we have 6 and 5 right? do the same with 4-4x = -8

OpenStudy (anonymous):

x=3

OpenStudy (loser66):

next is 4-4x= 0

OpenStudy (anonymous):

i got it thanks

OpenStudy (loser66):

ok